b/4+7=53....I know the steps but dont really understand how to solve. math handout
Answer by checkley75(3666)

B/4+7=53
B/7=53-7
B/4=46 NOW WE NEED TO CROSS MULTIPLY
B=46*4
B=184 ANSWER
PROOF
184/4+7=53
46+7=53
53=53